Twelve grams of nitrogen (N2) in a steel tank are hated from 25.0 to 125ºC. (a) Howmany moles of nitrogen are present? (b) How much hat is transferred to the nitrogen?

Grade:11

1 Answers

Navjyot Kalra
8 years ago
To find out the number of moles n of nitrogen, substitute 12 g for given mass m and 28 g/mol for M (molecular weight of nitrogen) in the equation n = m/M,
n = m/M
= (12 g)/(28 g/mol)
= 0.429 mol
Thus 0.429 moles of nitrogen are present.

Since nitrogen (N2) is a diatomic gas, therefore Cv = 5/2 R.
To obtain the heat Q which is transferred to the nitrogen, substitute 0.429 mol for n, 5/2 R for Cv and 125̊ C for Tf and 25̊ C for Ti in the equation Q = nCv(Tf –Ti),
Q = nCv(Tf –Ti)
= (0.429 mol) (5/2 R) (125̊ C-25̊ C)
= (0.429 mol) (5/2 ×8.31 J/mol. K) ((125 + 273) K-(25 + 273) K) (Since, R = 8.31 J/mol. K)
= (0.429 mol) (5/2 ×8.31 J/mol. K) (398 K-298 K)
= (0.429 mol) (5/2 ×8.31 J/mol. K) (100 K)
= 891 J
From the above observation we conclude that, the heat Q which is transferred to the nitrogen would be 891 J.